A monkey of mass 20kg rides on a 40 kg trolley moving with constant speed of 8m/s along a horizontal track. If the monkey jumps vertically to grab the overhanging branch of a tree, the speed of the trolley after the monkey has jumped off is-?

Grade:11

5 Answers

India VK
10 years ago

8 m/s

Consider just the trolley as the system. Since the monkey jumps vertically, it does not exert any horizontal force on the trolley, so the trolley''s momentum remains unchanged. So its speed remains the same.

Prakhar Tayal
10 years ago

12m/s

60*8=40*x

x=12m/s

Conservation of momentum!

Regards,

Prakhar

Vaibhav
10 years ago
mass of monkey m=20 kgmass of trolley M =40 kgspeed of trolley with monkey v =8m/sBy principle of conservation of momentum,(m+M)v=MVV=(m+M)v/M  =(20+40)8/40  =12 ans
rishabh
10 years ago

Is the answer equals to 8/3

Rashmi
5 years ago
Since the monkey jumped off the trolley in a vertical direction which only provides a vertical force.. And as we know there is no contribution of vertical force in horizontal direction so the momentum of the trolley remains same, so the speed also remains same as before i.e. 8 m/s.
Hope you people understand